OpenStudy (anonymous):

integral 1/sqrt(5-x^2+x) dx

OpenStudy (anonymous):

partial fractions or does the denominator not factor?

OpenStudy (anonymous):

\[\int\limits_{?}^{?} 1/\sqrt{5-x ^{2}+4x} dx \]

OpenStudy (anonymous):

no it doesn't factor. i guess you have to complete the damned square

OpenStudy (anonymous):

oh i didn't see the radical, sorry

OpenStudy (anonymous):

no intervals, just inegration. somebody please help me

OpenStudy (anonymous):

yeah, how would we start this

OpenStudy (anonymous):

complete the square on the inside of the radical, then arrange it to look like \[\frac{1}{\sqrt{a^2-u^2}}\] to give \(\sin^{-1}(u)\)or something close to it

OpenStudy (anonymous):

ohh ohh ohh okayy,, yeaaa i got it i got . thankyou...

OpenStudy (anonymous):

it is a colossal pain but doable i would cheat

OpenStudy (lgbasallote):

-5 + x^2 + 4? isn't that 9 - x^2? it's in the form already @satellite73

OpenStudy (anonymous):

no there is a 4x in the denominator

OpenStudy (anonymous):

yea it is 4x

OpenStudy (lgbasallote):

oh thought it was 4 ssorry :p dont mind me

OpenStudy (anonymous):

its alright